OpenStudy (anonymous):

Graph the two equations to solve the system. 3x + 2y = 6 y = -3/2 x - 1

OpenStudy (anonymous):

find the intersection between these two graphs

OpenStudy (anonymous):

Solve for y in the first equation first. Then graph and find the intersection as @higgs suggested.

OpenStudy (anonymous):

First move 3x to the right side by using inverse operation (subtraction). THEN get the y by itself through inverse operation (division)

OpenStudy (anonymous):

If you don't have a graphing calculator: http://my.hrw.com/math06_07/nsmedia/tools/Graph_Calculator/graphCalc.html Enter the two functions on the first page, then click graph on the right. Afterwards, if you can't tell where the intersection is located, click on the tab at the top labelled INTERSECTION, put check marks in the boxes next to the functions...and I think the rest is self-explanatory.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Let me know if you have a problem with it...

OpenStudy (anonymous):

wait how do i find out where the intersection is again

OpenStudy (anonymous):

it says there is no intersection

OpenStudy (anonymous):

hahaha i figured it out

OpenStudy (anonymous):

My fault!! I apologize for taking so long...glad you figured it out!! : )

OpenStudy (anonymous):

So you got two parallel lines right? Just to clarify.

OpenStudy (anonymous):

Parallel lines (lines that NEVER intersect) have the SAME slopes but DIFFERENT y-intercepts.

OpenStudy (anonymous):

Which is the case for this particular problem.

OpenStudy (anonymous):

no sorry i got stuck again! and its irritating me!

OpenStudy (anonymous):

Same problem or a new one? Sorry I was in the computer science forum.

OpenStudy (anonymous):

same problem but is it -5,14

OpenStudy (anonymous):

-14

OpenStudy (anonymous):

Cool...let's look at the first equation again...

OpenStudy (anonymous):

3x + 2y = 6

OpenStudy (anonymous):

We want to solve for y (i.e. get y alone)

OpenStudy (anonymous):

So we start by moving 3x to the right side of the equal sign. But we can only do that by using inverse (opposite) operations. Inverse operation: Since it's positive 3x (+3x), we move it using negative 3x (-3x)

OpenStudy (anonymous):

3x + 2y = 6 -3x + 2y = -3x + 6 ------------------- 0x + 2y = -3x + 6 2y = -3x + 6 Now we need to get the y by itself, again by using inverse (opposite) operation. Inverse operation: Since it's 2y (2 times y) and we want y alone, we need to divide (2y/2). 2y/2 = -3x/2 + 6/2 Simplifies too...

OpenStudy (anonymous):

to*...

OpenStudy (anonymous):

y = -3/2x + 3

OpenStudy (anonymous):

The 2s on the left cancel each other out. And 6/2 = 3 on the far right. The fraction next to x stays as it is.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The 2nd equation is: y = -3/2x -1

OpenStudy (anonymous):

They BOTH share the SAME slope, but have DIFFERENT y-intercepts (one y-interecept is 3, the other y-intercept is -1)

OpenStudy (anonymous):

That's a math rule: PARALLEL (//) lines and equations, have the EXACT SAME slope but different y-intercepts.

OpenStudy (anonymous):

// = parallel, btw.

OpenStudy (anonymous):

So these two equations: y = -3/2x + 3 and y = -3/2x - 1 will NEVER intersect. Therefore there is: No Solution.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Or an empty set: \[{\emptyset}\]

OpenStudy (anonymous):

Good job. Take care

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem. Don't forget to select a Best Response. Take care